OpenStudy (cenaida):

Solve the following system of equations using elimination: 13x + 5y = 16 2x - 8y = 20 A. (-2, -3) B. (-2, 2) C. (2, -2) D. (2, 2)

OpenStudy (john_es):

These systems grow like flowers in the field. :) You must multiply the first by 8 and the second by 5, \[\begin{cases}104x+40y=128\\ 10x-40y=100\end{cases}\]And sum both equations, \[114x=228\Rightarrow x=2\]Now you can susbtitute this value in the second equation and obtains for y, y=-2. So you have (2,-2)

OpenStudy (anonymous):

C